OpenStudy (anonymous):

The formula A=22.9e^0.0183t models the population of Texas, A, in millions, t years after 2005. a)-- b) when will the population of Texas reach 27 million?

OpenStudy (anonymous):

Plug in: 27 = 22.9e^(0.0183t) Divide by 22.9: 27/22.9 = e^(0.0183t) Take the natural logarithm: ln(27/22.9) = ln(e^(0.0183t)) ln(27/22.9) = 0.0183t Divide by 0.0183: ln(27/22.9)/0.0183 = t t ≈ 8.999 Would this be correct?

OpenStudy (campbell_st):

well substitute it and check..

OpenStudy (anonymous):

But wait idk how to find "What was the population of texas in 2005" im stuck there

OpenStudy (campbell_st):

its extremely close... of you had used t = 9 you get the correct answer

OpenStudy (anonymous):

dude but im serious idk how to do this; its the only problem i dont get>.<

jimthompson5910 (jim_thompson5910):

it's not 9, but more like 8.999997565 which I guess is close enough

OpenStudy (campbell_st):

well 2005 let t = 0 so\[A_{0} = 22 900 000\]

OpenStudy (campbell_st):

well I found \[22.9 \times e^{0.0183 \times 9} = 27\]

OpenStudy (campbell_st):

and Ao is the initial population

jimthompson5910 (jim_thompson5910):

rounding error maybe

OpenStudy (anonymous):

^ Wait for which one?

OpenStudy (campbell_st):

so 9 years after 2005 is...

OpenStudy (campbell_st):

thats when the population will reach 27 mill

jimthompson5910 (jim_thompson5910):

and the beauty of the general equation A = P*e^(kt) is that when t =0, the initial population is P So in this case, when t = 0 (ie the year 2005), the population is 22.9 million

OpenStudy (anonymous):

I get you campbell but i dont get you jim where do you plug in what?

jimthompson5910 (jim_thompson5910):

A = 22.9e^(0.0183t) A = 22.9e^(0.0183*0) ... plug in t = 0 A = 22.9e^(0) A = 22.9(1) A = 22.9

jimthompson5910 (jim_thompson5910):

So when t = 0, A = 22.9 Which means that in the year 2005, the population is 22.9 million

OpenStudy (anonymous):

Thats the population f texas in 2005 right?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thanks!:D

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (campbell_st):

glad to help

OpenStudy (anonymous):

Dude you both guys just help me big time! Im really thankful no lie!:D

jimthompson5910 (jim_thompson5910):

lol that's great that we were able to help you out